# Investigating Adversarial Attacks in Software Analytics via Machine Learning Explainability

### v1

@@ 3.5. Robustness Evaluation
4. Experimental Results and Analysis
We implement and evaluate our experiment using the scikit-learn (Buitinck et al., 2013), SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016), PyExplainer (Pornprasit et al., 2021), and Adversarial Robustness Toolbox (ART) (Nicolae et al., 2018) libraries. Our study involves six datasets, three ML explainability techniques, and seven classical ML models, resulting in 126 ($6\times 3\times 7=126$) experimental combinations. Similarly to prior studies (Yang et al., 2022; Hashemi and Fathi, 2020; Ballet et al., 2019; Zhang et al., 2020), we focus exclusively on the proportion of correctly predicted instances in the test dataset. Thus, we start by evaluating the performance of our chosen ML models, and then, we assess the effectiveness of our proposed approach through extensive experiments by addressing the following two research questions:
@@ 4. Experimental Results and Analysis
Baseline Attacks: To the best of our knowledge, adversarial attacks targeting classical ML models trained on tabular data for different software analytics tasks, such as JIT defect prediction, code review comment classification, and clone detection, have not been explored. Hence, we compare our approach against four state-of-the-art attack techniques: Zoo (Chen et al., 2017b), Boundary attack (Brendel et al., 2017), PermuteAttack (Hashemi and Fathi, 2020), and HopSkipJump (Chen et al., 2020), specifically applied for ML models trained on tabular data. We deliberately selected these as baselines due to the black-box nature of adversarial attacks and their publicly available implementations. We briefly describe each of the baseline attack approaches below:
@@
    (3) $\|x_{orig}-x_{cnt}\|_{0}\leq\delta_{0,max}\text{ and }\|x_{orig}-x_{cnt}\|_{2}\leq\delta_{2,max}$
    (4) $\text{ComputeFitness}(\text{X})=\|f(x)_{t}-f(x_{orig})_{t}\|_{2}-\rho_{0}\|x_{orig}-x\|_{0}-\rho_{1}\|x_{orig}-x\|_{2}$
@@
4.1. Analyse Model Performance
@@
Table 3. Optimized hyperparameters for different ML models on the CLCDSA dataset after tuning.
ML Model Hyperparameters
LR C: 1.0, dual: False, max_iter: 140, penalty: ‘l2’, solver: ‘lbfgs’
DT criterion: ‘entropy’, max_depth: 15, min_samples_leaf: 20, min_samples_split: 8
RF bootstrap: False, max_depth: None, max_features: ‘sqrt’, min_samples_leaf: 1, min_samples_split: 2, n_estimators: 50
MLP solver: ‘adam’, learning_rate: ‘adaptive’, hidden_layer_sizes: (10, 30, 10), alpha: 0.05, activation: ‘relu’
ADA algorithm: ‘SAMME.R’, learning_rate: 1.02, n_estimators: 20
BAG max_features: 13, max_samples: 100, n_estimators: 800
GBC learning_rate: 1, max_depth: 7, min_samples_leaf: 0.1, min_samples_split: 0.1, n_estimators: 200
Table 4 presents the accuracy, F1-score, and AUC values for our ML models trained on various datasets without adversarial attacks. Notably, when considering the cross-project mobile apps dataset, the LR model displays the lowest AUC value (0.73), while the BAG model achieves the highest AUC value (0.85). On the other hand, ML models trained on the CLCDSA dataset consistently exhibit high AUC values, ranging from 0.82 to 0.96. In contrast, for the code review dataset, Table 4 reveals relatively lower AUC values; however, the AUC values range from 0.63 to 0.96 for all the ML models. A closer look at Table 4 emphasizes that all the ML models maintain acceptable accuracy, F1-scores, and AUC values, indicating their high accuracy and non-overfitting characteristics. Furthermore, all test-AUCs (with the exception of two cases and the code review dataset) exceed 0.75, a benchmark recommended by previous studies as the minimum required to guarantee the reliability of explanations (Roy et al., 2022; Lyu et al., 2021).
Table 4. Performance of the ML models on different datasets without adversarial attacks considering different evaluation metrics such as Accuracy (Acc), F1-score (F1), and AUC values.
Dataset
Cross Project Java Project Postgres CLCDSA Code Review BigCloneBench
ML Model Acc F-1 AUC Acc F-1 AUC Acc F-1 AUC Acc F-1 AUC Acc F-1 AUC Acc F-1 AUC
LR 0.68 0.49 0.73 0.75 0.37 0.76 0.75 0.57 0.76 0.76 0.81 0.82 0.62 0.69 0.63 0.75 0.74 0.76
DT 0.79 0.60 0.83 0.80 0.49 0.80 0.78 0.59 0.79 0.86 0.88 0.93 0.67 0.73 0.68 0.84 0.84 0.85
RF 0.81 0.58 0.84 0.88 0.55 0.85 0.80 0.59 0.82 0.91 0.93 0.96 0.71 0.76 0.72 0.88 0.87 0.88
MLP 0.75 0.57 0.83 0.80 0.48 0.81 0.71 0.51 0.73 0.83 0.86 0.92 0.60 0.68 0.63 0.83 0.83 0.84
ADA 0.73 0.56 0.81 0.79 0.47 0.81 0.78 0.58 0.79 0.81 0.77 0.85 0.63 0.65 0.67 0.83 0.82 0.83
BAG 0.77 0.59 0.85 0.78 0.50 0.83 0.78 0.61 0.82 0.81 0.84 0.89 0.65 0.73 0.71 0.9 0.89 0.9
GBC 0.81 0.59 0.84 0.81 0.48 0.80 0.78 0.57 0.80 0.85 0.89 0.94 0.63 0.69 0.65 0.87 0.86 0.87
4.2. Present Experimental Results
@@
Figure 3. selecting the top-$k$ feature values for different datasets. The first column (e.g., 3(a) for cross-project mobile apps, 3(c) for Java projects, 3(e) for CLCDSA, and 3(g) for BCB) shows how the ASR metric value changes when different combinations of features are selected from the feature importance rank using SHAP explainability. The second column (e.g., 3(b) for cross-project mobile apps, 3(d) for Java projects, 3(f) for CLCDSA, and 3(h) for BCB) shows how the ASR metric value changes when different combinations of features are selected from the feature importance rank using LIME explainability.
@@
Figures 4 and 5 illustrate the distribution of prediction probability differences between the original and transformed instances for the ML models trained on the cross-project mobile apps and BCB datasets. The both figures clearly indicate substantial differences in prediction probabilities between the original and transformed instances, except for the ADA model and SHAP (Figures 4(e) and 5(e)). For example, in the case of the LR model, the mean difference is approximately $0.61$ when changing the top-$k$ important feature values identified by SHAP and LIME. The mean probability difference is approximately $0.47$ when modifying the top-$k$ important feature values identified by PyExplainer (PyExp). The mean difference consistently exceeds $0.22$ for all other ML models and explainability techniques. This denotes a significant prediction probability difference exists between the original and the transformed instances when we modify the original instance based on the top-$k$ important features identified by ML explainability techniques. The only exception is in the case of the ADA model, where SHAP shows a comparatively lower prediction probability difference. In contrast, LIME and PyExplainer maintain consistency compared to SHAP for all the ML models.
@@
Figure 4. Distribution of prediction probability differences between the original and transformed instances for various ML models trained on the cross-project mobile apps dataset are shown in Figures 4(a), 4(b), 4(c), 4(d), 4(e), 4(f), and 4(g). These figures represent the distribution for LR, DT, RF, MLP, ADA, BAG, and GBC models, respectively.
@@
Figure 5. Distribution of prediction probability differences between the original and transformed instances for various ML models trained on the BigCloneBench dataset are shown in Figures 5(a), 5(b), 5(c), 5(d), 5(e), 5(f), and 5(g). These figures represent the distribution for LR, DT, RF, MLP, ADA, BAG, and GBC models, respectively.
@@
Table 5 presents the ASR metric and top-$k$ values both for the explanation-guided adversarial attacks and the selected state-of-the-art adversarial attacks on tabular data. The ASR metric quantifies the degree to which the models’ accuracy is compromised when specific features are altered to generate adversarial examples. For instance, in Table 5, it is evident that the BAG model trained on the cross-project mobile apps dataset fails to accurately predict $76.4\%$ of instances that it correctly predicted before adversarial attacks when considering SHAP. Similarly, the LR model fails to predict $75.6\%$ and $56.3\%$ instances accurately after adversarial attacks when considering LIME and PyExplainer, respectively. The ASR metric value for the cross-project mobile apps dataset ranges between $23.3\%$ and $76.4\%$. Similarly, the range is 14.8%–53.6% for the Java project dataset, 11.5%–86.6% for the Postgres dataset, 13.4%–61.4% for the CLCDSA dataset, 17.87%–69.0% for the BCB dataset, and 26.9%–71.6% for the code review dataset, respectively. In summary, the ASR metric value ranges between 11.5%–86.6% for all the ML models, considering all the datasets.
Table 5. Attack Success Rate (ASR) metric values and the Top-$k$ modified features (e.g., a minimal $\ell_{0}$ perturbation) when we apply explanation-guided adversarial attacks and selected baseline attacks on the ML models for various datasets.
Cross-project Mobile Apps
SHAP LIME PyExplainer PermuteAttack Zoo Boundary HopSkipJump
ML Model ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$
LR 75.8 Top-1 75.6 Top-1 56.3 88.1 Top-4 78.2 Top-6 56.8 Top-6 55.7 Top-6
DT 58.4 Top-2 58.2 Top-2 23.3 99.6 Top-3 25.2 Top-2 85.7 Top-6 84.9 Top-6
RF 60.8 Top-3 48.2 Top-2 32.1 98.1 Top-4 40.8 Top-3 99.7 Top-6 99.9 Top-6
MLP 70.7 Top-3 54.1 Top-2 38.7 92.1 Top-4 71.5 Top-6 99.9 Top-6 99.9 Top-6
ADA 72.2 Top-2 57.7 Top-2 48.0 11.4 Top-2 5.1 Top-2 23.3 Top-6 23.1 Top-6
BAG 76.4 Top-2 57.7 Top-2 28.6 74.1 Top-4 31.1 Top-6 22.2 Top-6 22.2 Top-6
GBC 71.4 Top-3 47.1 Top-2 28.4 99.3 Top-4 42.8 Top-2 84.4 Top-6 84.9 Top-6
Java Project
LR 51.5 Top-6 32.9 Top-2 32.1 99.4 Top-4 35.5 Top-13 11.4 Top-27 11.8 Top-27
DT 53.5 Top-2 45.2 Top-1 18.4 98.4 Top-5 0.0 Top-0 87.9 Top-27 87.9 Top-27
RF 29.1 Top-2 27.3 Top-1 14.8 98.6 Top-7 0.0 Top-0 99.9 Top-27 100 Top-27
MLP 53.5 Top-1 44.3 Top-1 26.3 99.6 Top-6 26.1 Top-13 14.2 Top-27 14.3 Top-27
ADA 53.6 Top-3 22.4 Top-2 32.9 0.0 Top-0 0.0 Top-0 12.4 Top-27 12.4 Top-27
BAG 38.3 Top-2 35.8 Top-1 21.7 73.1 Top-7 0.25 Top-3 99.7 Top-27 99.8 Top-27
GBC 45.4 Top-2 38.6 Top-1 22.2 99.6 Top-6 0.0 Top-0 11.5 Top-27 11.4 Top-27
Postgres
LR 86.6 Top-1 83.4 Top-2 32.0 99.4 Top-5 99.5 Top-9 19.9 Top-12 19.9 Top-12
DT 80.3 Top-2 82.8 Top-2 25.0 100 Top-4 88.7 Top-3 35.9 Top-12 33.2 Top-12
RF 70.9 Top-2 71.6 Top-3 20.7 99.6 Top-5 91.6 Top-4 45.5 Top-12 45.3 Top-12
MLP 67.2 Top-1 63.9 Top-3 36.6 93.5 Top-6 91.1 Top-12 31.9 Top-12 30.8 Top-12
ADA 77.3 Top-1 64.4 Top-1 21.7 5.6 Top-2 66.8 Top-3 19.9 Top-12 19.9 Top-12
BAG 66.7 Top-2 72.3 Top-2 11.5 100 Top-4 87.2 Top-7 29.4 Top-12 29.4 Top-12
GBC 57.3 Top-2 63.7 Top-2 30.7 100 Top-5 88.2 Top-4 88.2 Top-12 90.1 Top-12
CLCDSA
LR 34.5 Top-1 35.5 Top-1 55.2 99.0 Top-6 9.7 Top-13 100 Top-18 100 Top-18
DT 53.9 Top-2 40.9 Top-2 13.4 95.5 Top-5 0.0 Top-0 74.8 Top-18 74.5 Top-18
RF 46.1 Top-2 36.5 Top-2 15.3 96.2 Top-8 0.0 Top-0 64.1 Top-18 63.5 Top-18
MLP 61.4 Top-2 49.8 Top-3 36.8 96.3 Top-6 15.3 Top-13 100 Top-18 100 Top-18
ADA 51.1 Top-2 42.1 Top-2 17.3 0.0 Top-0 0.0 Top-0 96.2 Top-18 96.1 Top-18
BAG 28.1 Top-3 40.2 Top-2 16.1 82.9 Top-5 0.63 Top-7 84.2 Top-18 85.2 Top-18
GBC 34.3 Top-2 33.3 Top-2 14.9 98.5 Top-6 0.0 Top-0 88.3 Top-18 87.8 Top-18
Code Review
LR 53.9 Top-1 48.9 Top-1 41.9 99.3 Top-3 100 Top-10 100 Top-15 100 Top-15
DT 38.0 Top-2 35.3 Top-3 26.9 99.6 Top-3 50.0 Top-2 68.0 Top-15 66.0 Top-15
RF 39.3 Top-3 32.9 Top-2 21.1 99.3 Top-4 86.7 Top-4 66.5 Top-15 66.4 Top-15
MLP 71.6 Top-1 57.5 Top-2 28.3 79.8 Top-5 89.5 Top-13 100 Top-15 100 Top-15
ADA 62.9 Top-1 47.9 Top-2 38.1 40.0 Top-3 66.4 Top-3 57.8 Top-15 52.1 Top-15
BAG 27.6 Top-2 23.5 Top-2 26.9 79.3 Top-4 80.7 Top-8 100 Top-15 100 Top-15
GBC 44.4 Top-2 38.1 Top-3 26.9 100 Top-4 70.4 Top-4 92.9 Top-15 89.4 Top-15
BigCloneBench
LR 68.0 Top-1 56.2 Top-1 52.4 100 Top-3 100 Top-5 99.9 Top-6 99.9 Top-6
DT 50.9 Top-1 63.5 Top-3 17.8 100 Top-3 87.8 Top-4 51.2 Top-6 51.2 Top-6
RF 53.8 Top-2 64.9 Top-2 20.5 99.9 Top-5 98.6 Top-5 52.6 Top-6 52.6 Top-6
MLP 69.5 Top-1 68.6 Top-2 44.4 100 Top-4 100 Top-6 99.4 Top-6 99.0 Top-6
ADA 69.0 Top-1 68.9 Top-2 37.3 92.6 Top-3 99.6 Top-5 53.3 Top-6 53.3 Top-6
BAG 64.7 Top-1 63.0 Top-2 17.8 100 Top-5 97.2 Top-5 53.0 Top-6 53.0 Top-6
GBC 67.6 Top-1 67.8 Top-2 18.7 99.8 Top-5 95.5 Top-5 52.2 Top-6 52.2 Top-6
@@
Table 6. Performance of the ML models on different datasets without adversarial attacks considering different evaluation metrics such as Accuracy (Acc), F1-score (F1), and AUC values before balancing (BB) and after balancing (AB) the test data.
@@
Table 7. Attack Success Rate (ASR) metric values and the Top-$k$ modified features (e.g., a minimal $\ell_{0}$ perturbation) when we apply explanation-guided adversarial attacks and selected baseline attacks on the ML models for the balanced Java project and Postgres datasets.
Java Project
SHAP LIME PyExplainer PermuteAttack Zoo Boundary HopSkipJump
ML Model ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$
LR 40.98 Top-3 48.78 Top-3 25.06 97.98 Top-7 23.38 Top-13 45.43 Top-27 44.98 Top-27
DT 60.71 Top-2 47.49 Top-1 20.59 97.95 Top-5 0.0 Top-0 55.12 Top-27 55.12 Top-27
RF 44.58 Top-2 39.58 Top-3 9.07 98.75 Top-6 0.0 Top-0 100 Top-27 100 Top-27
MLP 50.09 Top-2 38.4 Top-1 22.26 99.27 Top-7 16.67 Top-12 45.82 Top-27 45.08 Top-27
ADA 46.62 Top-3 41.59 Top-3 19.11 0.0 Top-0 0.0 Top-0 46.61 Top-27 46.42 Top-27
BAG 51.96 Top-2 44.21 Top-2 17.86 56.19 Top-6 0.02 Top-1 99.63 Top-27 99.63 Top-27
GBC 55.97 Top-3 46.27 Top-1 14.37 99.25 Top-6 0.0 Top-0 41.6 Top-27 41.6 Top-27
Postgres
LR 85.89 Top-1 83.06 Top-3 34.54 99.46 Top-6 99.32 Top-8 41.8 Top-12 41.8 Top-12
DT 85.63 Top-2 82.92 Top-2 6.9 85.87 Top-4 85.87 Top-12 66.54 Top-12 68.55 Top-12
RF 65.45 Top-2 72.91 Top-2 21.36 100 Top-4 91.99 Top-4 86.08 Top-12 84.77 Top-12
MLP 56.99 Top-1 59.45 Top-2 22.4 87.94 Top-6 88.08 Top-11 48.9 Top-12 48.76 Top-12
ADA 84.09 Top-3 56.98 Top-1 19.61 8.34 Top-3 56.97 Top-4 40.93 Top-12 40.93 Top-12
BAG 83.59 Top-2 79.56 Top-3 36.98 100 Top-4 86.71 Top-7 50.13 Top-12 49.47 Top-12
GBC 70.99 Top-2 73.54 Top-2 19.75 100 Top-4 85.69 Top-4 93.24 Top-12 92.05 Top-12

### v2

@@ 3.5 Robustness Evaluation
4 Experimental Results and Analysis
We implement and evaluate our experiment using the scikit-learn [83], SHAP [57], LIME [58], PyExplainer [35], and Adversarial Robustness Toolbox (ART) [84] libraries. Our study involves six datasets, three ML explainability techniques, and seven classical ML models, resulting in 126 ($6\times 3\times 7=126$) experimental combinations. Similarly to prior studies [25, 61, 80, 27], we focus exclusively on the proportion of correctly predicted instances in the test dataset. Thus, we start by evaluating the performance of our chosen ML models, and then, we assess the effectiveness of our proposed approach through extensive experiments by addressing the following two research questions:
@@ 4 Experimental Results and Analysis
Baseline Attacks: To the best of our knowledge, adversarial attacks targeting classical ML models trained on tabular data for various software analytics tasks, such as JIT defect prediction, code review comment classification, and clone detection, have not been explored. Since our proposed attack technique, similar to existing studies [81, 76, 31, 85], operates under a black-box setting (i.e., it does not require access to model internals such as gradients or parameters), we compare it against four state-of-the-art black-box attack techniques: Zoo [59], Boundary Attack [60], PermuteAttack [61], and HopSkipJump [62], all specifically designed for ML models trained on tabular data. We deliberately selected these baselines due to their black-box nature and the availability of their public implementations. A brief description of each baseline attack technique is provided below.
@@
    $\|x_{orig}-x_{cnt}\|_{0}\leq\delta_{0,max}\text{ and }\|x_{orig}-x_{cnt}\|_{2}\leq\delta_{2,max}$ (3)
    $\text{ComputeFitness}(\text{X})=\|f(x)_{t}-f(x_{orig})_{t}\|_{2}-\rho_{0}\|x_{orig}-x\|_{0}-\rho_{1}\|x_{orig}-x\|_{2}$ (4)
@@
4.1 Analyse Model Performance
@@
Table 4: Optimized hyperparameters for different ML models on the CLCDSA dataset after tuning
ML Model Hyperparameters LR C: 1.0, dual: False, max_iter: 140, penalty: ‘l2’, solver: ‘lbfgs’ DT criterion: ‘entropy’, max_depth: 15, min_samples_leaf: 20, min_samples_split: 8 RF bootstrap: False, max_depth: None, max_features: ‘sqrt’, min_samples_leaf: 1, min_samples_split: 2, n_estimators: 50 MLP solver: ‘adam’, learning_rate: ‘adaptive’, hidden_layer_sizes: (10, 30, 10), alpha: 0.05, activation: ‘relu’ ADA algorithm: ‘SAMME.R’, learning_rate: 1.02, n_estimators: 20 BAG max_features: 13, max_samples: 100, n_estimators: 800 GBC learning_rate: 1, max_depth: 7, min_samples_leaf: 0.1, min_samples_split: 0.1, n_estimators: 200
Table 5 presents the accuracy, F1-score, and AUC values for our ML models trained on various datasets without adversarial attacks. Notably, when considering the cross-project mobile apps dataset, the LR model displays the lowest AUC value (0.73), while the BAG model achieves the highest AUC value (0.85). On the other hand, ML models trained on the CLCDSA dataset consistently exhibit high AUC values, ranging from 0.82 to 0.96. In contrast, for the code review dataset, Table 5 reveals relatively lower AUC values; however, the AUC values range from 0.63 to 0.96 for all the ML models. A closer look at Table 5 emphasizes that all the ML models maintain acceptable accuracy, F1-scores, and AUC values, indicating their high accuracy and non-overfitting characteristics. Furthermore, all test-AUCs (with the exception of two cases and the code review dataset) exceed 0.75, a benchmark recommended by previous studies as the minimum required to guarantee the reliability of explanations [65, 75].
Table 5: Performance of the ML models on different datasets without adversarial attacks considering different evaluation metrics such as Accuracy (Acc), F1-score (F1), and AUC values
Dataset Cross Project Java Project Postgres CLCDSA Code Review BigCloneBench ML Model Acc F-1 AUC Acc F-1 AUC Acc F-1 AUC Acc F-1 AUC Acc F-1 AUC Acc F-1 AUC LR 0.68 0.49 0.73 0.75 0.37 0.76 0.75 0.57 0.76 0.76 0.81 0.82 0.62 0.69 0.63 0.75 0.74 0.76 DT 0.79 0.60 0.83 0.80 0.49 0.80 0.78 0.59 0.79 0.86 0.88 0.93 0.67 0.73 0.68 0.84 0.84 0.85 RF 0.81 0.58 0.84 0.88 0.55 0.85 0.80 0.59 0.82 0.91 0.93 0.96 0.71 0.76 0.72 0.88 0.87 0.88 MLP 0.75 0.57 0.83 0.80 0.48 0.81 0.71 0.51 0.73 0.83 0.86 0.92 0.60 0.68 0.63 0.83 0.83 0.84 ADA 0.73 0.56 0.81 0.79 0.47 0.81 0.78 0.58 0.79 0.81 0.77 0.85 0.63 0.65 0.67 0.83 0.82 0.83 BAG 0.77 0.59 0.85 0.78 0.50 0.83 0.78 0.61 0.82 0.81 0.84 0.89 0.65 0.73 0.71 0.9 0.89 0.9 GBC 0.81 0.59 0.84 0.81 0.48 0.80 0.78 0.57 0.80 0.85 0.89 0.94 0.63 0.69 0.65 0.87 0.86 0.87
4.2 Present Experimental Results
@@
Figure 3: selecting the top-$k$ feature values for different datasets. The first column (e.g., 3(a) for cross-project mobile apps, 3(c) for Java projects, 3(e) for CLCDSA, and 3(g) for BCB) shows how the ASR metric value changes when different combinations of features are selected from the feature importance rank using SHAP explainability. The second column (e.g., 3(b) for cross-project mobile apps, 3(d) for Java projects, 3(f) for CLCDSA, and 3(h) for BCB) shows how the ASR metric value changes when different combinations of features are selected from the feature importance rank using LIME explainability.
@@
Figures 4 and 5 illustrate the distribution of prediction probability differences between the original and transformed instances for the ML models trained on the cross-project mobile apps and BCB datasets. The both figures clearly indicate substantial differences in prediction probabilities between the original and transformed instances, except for the ADA model and SHAP (Figures 4(e) and 5(e)). For example, in the case of the LR model, the mean difference is approximately $0.61$ when changing the top-$k$ important feature values identified by SHAP and LIME. The mean probability difference is approximately $0.47$ when modifying the top-$k$ important feature values identified by PyExplainer (PyExp). The mean difference consistently exceeds $0.22$ for all other ML models and explainability techniques. This denotes that a significant prediction probability difference exists between the original and the transformed instances when we modify the original instance based on the top-$k$ important features identified by ML explainability techniques. The only exception is in the case of the ADA model, where SHAP shows a comparatively lower prediction probability difference. In contrast, LIME and PyExplainer maintain consistency compared to SHAP for all the ML models.
@@
Figure 4: Distribution of prediction probability differences between the original and transformed instances for various ML models trained on the cross-project mobile apps dataset are shown in Figures 4(a), 4(b), 4(c), 4(d), 4(e), 4(f), and 4(g). These figures represent the distribution for LR, DT, RF, MLP, ADA, BAG, and GBC models, respectively.
@@
Figure 5: Distribution of prediction probability differences between the original and transformed instances for various ML models trained on the BigCloneBench dataset are shown in Figures 5(a), 5(b), 5(c), 5(d), 5(e), 5(f), and 5(g). These figures represent the distribution for LR, DT, RF, MLP, ADA, BAG, and GBC models, respectively.
While Figures 4 and 5 visually demonstrate the differences in prediction probabilities between the original and transformed instances, we further perform a Wilcoxon Signed-Rank test [86] to assess these differences statistically. We also report Cliff’s $|\delta|$ effect size [87], whose interpretation is provided in Table 6. Our results show that, in all cases, the $p$-values are below the significance threshold (0.05), and the effect sizes exceed 0.74, except for the ADA model trained on the cross-project mobile apps and BigCloneBench datasets using the SHAP and LIME explainers. These findings provide strong statistical evidence of a significant difference in prediction probabilities between the original and transformed instances.
Table 6: Interpretation of Cliff’s $|\delta|$ effect size
Interval Interpret Effect
$\delta<0.147$ Negligible effect
$0.147\leq 0.33$ Small effect
$0.33\leq 0.474$ Medium effect
$\delta\geq 0.474$ Large effect
@@
Table 7: Attack Success Rate (ASR) metric values and the Top-$k$ modified features (e.g., a minimal $\ell_{0}$ perturbation) when we apply explanation-guided adversarial attacks and selected baseline attacks on the ML models for various datasets
Cross-project Mobile Apps SHAP LIME PyExplainer PermuteAttack Zoo Boundary HopSkipJump ML Model ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ LR 75.8 Top-1 75.6 Top-1 56.3 88.1 Top-4 78.2 Top-6 56.8 Top-6 55.7 Top-6 DT 58.4 Top-2 58.2 Top-2 23.3 99.6 Top-3 25.2 Top-2 85.7 Top-6 84.9 Top-6 RF 60.8 Top-3 48.2 Top-2 32.1 98.1 Top-4 40.8 Top-3 99.7 Top-6 99.9 Top-6 MLP 70.7 Top-3 54.1 Top-2 38.7 92.1 Top-4 71.5 Top-6 99.9 Top-6 99.9 Top-6 ADA 72.2 Top-2 57.7 Top-2 48.0 11.4 Top-2 5.1 Top-2 23.3 Top-6 23.1 Top-6 BAG 76.4 Top-2 57.7 Top-2 28.6 74.1 Top-4 31.1 Top-6 22.2 Top-6 22.2 Top-6 GBC 71.4 Top-3 47.1 Top-2 28.4 99.3 Top-4 42.8 Top-2 84.4 Top-6 84.9 Top-6 Java Project LR 51.5 Top-6 32.9 Top-2 32.1 99.4 Top-4 35.5 Top-13 11.4 Top-27 11.8 Top-27 DT 53.5 Top-2 45.2 Top-1 18.4 98.4 Top-5 0.0 Top-0 87.9 Top-27 87.9 Top-27 RF 29.1 Top-2 27.3 Top-1 14.8 98.6 Top-7 0.0 Top-0 99.9 Top-27 100 Top-27 MLP 53.5 Top-1 44.3 Top-1 26.3 99.6 Top-6 26.1 Top-13 14.2 Top-27 14.3 Top-27 ADA 53.6 Top-3 22.4 Top-2 32.9 0.0 Top-0 0.0 Top-0 12.4 Top-27 12.4 Top-27 BAG 38.3 Top-2 35.8 Top-1 21.7 73.1 Top-7 0.25 Top-3 99.7 Top-27 99.8 Top-27 GBC 45.4 Top-2 38.6 Top-1 22.2 99.6 Top-6 0.0 Top-0 11.5 Top-27 11.4 Top-27 Postgres LR 86.6 Top-1 83.4 Top-2 32.0 99.4 Top-5 99.5 Top-9 19.9 Top-12 19.9 Top-12 DT 80.3 Top-2 82.8 Top-2 25.0 100 Top-4 88.7 Top-3 35.9 Top-12 33.2 Top-12 RF 70.9 Top-2 71.6 Top-3 20.7 99.6 Top-5 91.6 Top-4 45.5 Top-12 45.3 Top-12 MLP 67.2 Top-1 63.9 Top-3 36.6 93.5 Top-6 91.1 Top-12 31.9 Top-12 30.8 Top-12 ADA 77.3 Top-1 64.4 Top-1 21.7 5.6 Top-2 66.8 Top-3 19.9 Top-12 19.9 Top-12 BAG 66.7 Top-2 72.3 Top-2 11.5 100 Top-4 87.2 Top-7 29.4 Top-12 29.4 Top-12 GBC 57.3 Top-2 63.7 Top-2 30.7 100 Top-5 88.2 Top-4 88.2 Top-12 90.1 Top-12 CLCDSA LR 34.5 Top-1 35.5 Top-1 55.2 99.0 Top-6 9.7 Top-13 100 Top-18 100 Top-18 DT 53.9 Top-2 40.9 Top-2 13.4 95.5 Top-5 0.0 Top-0 74.8 Top-18 74.5 Top-18 RF 46.1 Top-2 36.5 Top-2 15.3 96.2 Top-8 0.0 Top-0 64.1 Top-18 63.5 Top-18 MLP 61.4 Top-2 49.8 Top-3 36.8 96.3 Top-6 15.3 Top-13 100 Top-18 100 Top-18 ADA 51.1 Top-2 42.1 Top-2 17.3 0.0 Top-0 0.0 Top-0 96.2 Top-18 96.1 Top-18 BAG 28.1 Top-3 40.2 Top-2 16.1 82.9 Top-5 0.63 Top-7 84.2 Top-18 85.2 Top-18 GBC 34.3 Top-2 33.3 Top-2 14.9 98.5 Top-6 0.0 Top-0 88.3 Top-18 87.8 Top-18 Code Review LR 53.9 Top-1 48.9 Top-1 41.9 99.3 Top-3 100 Top-10 100 Top-15 100 Top-15 DT 38.0 Top-2 35.3 Top-3 26.9 99.6 Top-3 50.0 Top-2 68.0 Top-15 66.0 Top-15 RF 39.3 Top-3 32.9 Top-2 21.1 99.3 Top-4 86.7 Top-4 66.5 Top-15 66.4 Top-15 MLP 71.6 Top-1 57.5 Top-2 28.3 79.8 Top-5 89.5 Top-13 100 Top-15 100 Top-15 ADA 62.9 Top-1 47.9 Top-2 38.1 40.0 Top-3 66.4 Top-3 57.8 Top-15 52.1 Top-15 BAG 27.6 Top-2 23.5 Top-2 26.9 79.3 Top-4 80.7 Top-8 100 Top-15 100 Top-15 GBC 44.4 Top-2 38.1 Top-3 26.9 100 Top-4 70.4 Top-4 92.9 Top-15 89.4 Top-15 BigCloneBench LR 68.0 Top-1 56.2 Top-1 52.4 100 Top-3 100 Top-5 99.9 Top-6 99.9 Top-6 DT 50.9 Top-1 63.5 Top-3 17.8 100 Top-3 87.8 Top-4 51.2 Top-6 51.2 Top-6 RF 53.8 Top-2 64.9 Top-2 20.5 99.9 Top-5 98.6 Top-5 52.6 Top-6 52.6 Top-6 MLP 69.5 Top-1 68.6 Top-2 44.4 100 Top-4 100 Top-6 99.4 Top-6 99.0 Top-6 ADA 69.0 Top-1 68.9 Top-2 37.3 92.6 Top-3 99.6 Top-5 53.3 Top-6 53.3 Top-6 BAG 64.7 Top-1 63.0 Top-2 17.8 100 Top-5 97.2 Top-5 53.0 Top-6 53.0 Top-6 GBC 67.6 Top-1 67.8 Top-2 18.7 99.8 Top-5 95.5 Top-5 52.2 Top-6 52.2 Top-6
Note: Bolded cells indicate cases where the attack techniques failed to produce any successful adversarial examples
@@
Table 7 presents the ASR metric and top-$k$ values both for the explanation-guided adversarial attacks and the selected state-of-the-art adversarial attacks on tabular data. The bolded cells highlighted in grey represent cases where the attack techniques failed to produce any successful adversarial examples. The ASR metric quantifies the degree to which the models’ accuracy is compromised when specific features are altered to generate adversarial examples. For instance, in Table 7, it is evident that the BAG model trained on the cross-project mobile apps dataset fails to accurately predict $76.4\%$ of instances that it correctly predicted before adversarial attacks when considering SHAP. Similarly, the LR model fails to predict $75.6\%$ and $56.3\%$ instances accurately after adversarial attacks when considering LIME and PyExplainer, respectively. The ASR metric value for the cross-project mobile apps dataset ranges between $23.3\%$ and $76.4\%$. Similarly, the range is 14.8%–53.6% for the Java project dataset, 11.5%–86.6% for the Postgres dataset, 13.4%–61.4% for the CLCDSA dataset, 17.87%–69.0% for the BCB dataset, and 26.9%–71.6% for the code review dataset, respectively. In summary, the ASR metric value ranges between 11.5%–86.6% for all the ML models, considering all the datasets.
@@
Table 8: Performance of the ML models on different datasets without adversarial attacks considering different evaluation metrics such as Accuracy (Acc), F1-score (F1), and AUC values before balancing (BB) and after balancing (AB) the test data